If there is a right triangle ABC and AB is 6 cm, angle A is 15 degrees, angle B is 90 degrees, and angle is 75 degrees? What is the area in simplest radical form?
Answer by nerdybill(5410)

If there is a right triangle ABC and AB is 6 cm, angle A is 15 degrees, angle B is 90 degrees, and angle is 75 degrees? What is the area in simplest radical form?
.
Draw a diagram of the triangle and label the sides and angles.
.
Applying trig, you would use:
tan = opposite/adjacent
tan(15) = height/6
height = 6tan(15)
height = 6(2-sqrt(3))
.
area of a triangle = 1/2 * height * base
area of a triangle = 1/2 * 6(2-sqrt(3)) * 6
area of a triangle = 3(2-sqrt(3)) * 6
area of a triangle = 18(2-sqrt(3))